A girl throws a rock horizontally with a speed of 12 m/s from a bridge. It falls 2.28 s before hitting the water below. Neglect air resistance.
(a) How high is the bridge from the water below?
(b) How far horizontally does the rock travel before striking the water?
Answer #1

a) a = 9.81 m/s^2
t = 2.28 s

d = ?
d = 1/2at^2 = 0.5*9.81*2.28^2 = 25.5 m

b) distance = v*t = 12*2.28 = 27.36 m